An object accelerates at 2.2 m/s2 for 3.0 seconds. If the final velocity is 15m/s what was the initial velocity?


1
Expert's answer
2022-01-18T09:35:43-0500

"v_i=v_f-at=15-2.2·3=8.4\text{ m/s}."